Stock As stock has a beta of 1.5, and its required return is 12.00%. Stock Bs beta is 0.80. If the risk-free rate is 4.75%, what is the required rate of return on Bs stock? (Hint: First find the market risk premium using information about stock A.) A. 7.97% O B. 8.62% ○ C. 8.98% ○ D, 9.21% O E. 9.58%
Answer #1

Under Capital Asset Pricing Model, Expected return = Risk free rate + beta*market risk premium

Using this formula and data for Stock A, first we need to calculate the market risk premium.

Expected return (Stock A) = Risk free rate + beta*market risk premium

0.12 = 0.0475 + 1.5*market risk premium

0.0725 = 1.5*market risk premium

Market risk premium = 0.0725/1.5

Market risk premium = 0.0483 = 4.83%

Expected return (Stock B) = Risk free rate + beta*market risk premium

Expected return (Stock B) = 0.0475 + 0.8*0.0483 = 0.0862

Expected return (Stock B) = 8.62%

Correct answer = B. 8.62%

Please give me your valuable feedback if the answer is helpful.